OpenStudy (vixella1):

10x - 10y = 1 x = y - 3 Solve the system of equations by substitution. A. (-29/20, 31/20) B. (1/2, 2/5) C. no solution D. coincident

OpenStudy (supernova_sonntag):

You might find more help in the mathematics section.

OpenStudy (studyloverhojea):

10x - 10y = 1 x = y - 3 put in y - 3 for x.

OpenStudy (studyloverhojea):

10(y-3) - 10y. Do distributive property. 10y - 30 - 10y = 1 subtract like terms which are 10y in this case. -30 does not equal 1 so the answer is no solutions.

alones (alones):

@Vixella1 hey there (; Just to remind, this belongs to math section, NOT english..